OpenStudy (anonymous):

Find the derivative of f directly from the limit definition of the derivative. f(x)=3/x

OpenStudy (e.mccormick):

When they say to do it that way, you will have to set up the limit of a difference quotient or something along those lines.

OpenStudy (luigi0210):

\[\frac{ f(a-h)-f(a) }{ h }\] Right?

OpenStudy (e.mccormick):

As h goes to 0.

OpenStudy (luigi0210):

\[\lim_{h \rightarrow 0}\frac{ f(a-h)-f(a) }{ h }\]

OpenStudy (e.mccormick):

Yep. That is one of the two ways to state the limit definition.

OpenStudy (luigi0210):

Plug them in and solve bro

OpenStudy (anonymous):

I just don't really understand how to use the difference quotient. I didn't understand it in algebra either :(

OpenStudy (luigi0210):

It probably wasn't a good idea to close the question then

OpenStudy (luigi0210):

and all it is really is just plugging stuff in and solving

OpenStudy (e.mccormick):

Remember function composition? \((f\circ g)(x)\) where you put one function into another? This is the same thing. You put \((a-h)\) into \(f(x)\) and the result goes where \(f(a-h)\) is.

OpenStudy (luigi0210):

we'll walk you though it if you want.. we got mccormick here :P

OpenStudy (luigi0210):

But yea just replace the x with a's and a-h's where needed. It's just like mccormick said

OpenStudy (anonymous):

That would be great

OpenStudy (anonymous):

There is only one x. maybe I am confused by simplicity? That is what I am hoping for.

OpenStudy (e.mccormick):

It does not matter if it s 1 x or 50. It gets the same process.

OpenStudy (e.mccormick):

It does not matter if it s 1 x or 50. It gets the same process. If I say to evaluate \(x=6\), you do \(f(6)=\frac{3}{6}\implies f(6)=\frac{1}{2}\). Well, this time we want to evaluate \((a-h)\) so \(f(a-h)=\) what? Once you have that, you put it into te right place on the difference quotent. Then you do the same thing with just a. Then you have fun solving it!

OpenStudy (e.mccormick):

hmm... some of that got repeated by accident. LOL

OpenStudy (anonymous):

LOL I just did it too

OpenStudy (anonymous):

I was trying to say maybe you could show me how you would find the derivative of 5/x using the limit definition? I am having a tough time computing that into what I need to do.

OpenStudy (anonymous):

I may have just figured it out

OpenStudy (luigi0210):

You got this bro!

OpenStudy (e.mccormick):

The limit definition of a derivative is stated one of two ways: \[\lim_{h \rightarrow 0}\frac{ f(a-h)-f(a) }{ h }\]or \[\lim_{x \rightarrow a}\frac{ f(x)-f(a) }{ x-a }\]Usually the first is easier.

OpenStudy (luigi0210):

Isn't the second one used for finding it at a certain point?

OpenStudy (e.mccormick):

It can be used for when it becomes 0 as well. In fact, the first comes from the second.

OpenStudy (anonymous):

So: \[\frac{ \frac{ 3 }{ a-h }-\frac{ 3 }{ a } }{ h }\]? Then slove from here?

OpenStudy (e.mccormick):

ALMOST! You need to keep the limit part on there until you evaluate the limit.

OpenStudy (anonymous):

Perfct, thank you!

OpenStudy (luigi0210):

But yea you're getting it, good job!

OpenStudy (luigi0210):

And if you did it right the answer should be -3/x^2

OpenStudy (luigi0210):

or -3/a^2

OpenStudy (e.mccormick):

Yah, all the algebra to get it where:\[\lim_{h \rightarrow 0}\frac{ \frac{ 3 }{ a-h }-\frac{ 3 }{ a } }{ h }\]won't have 0 for a denominator, then it should simplify to something like what Luigi0210 is showing. When they say "limit definition" the goal is to have you show all the work between.

OpenStudy (luigi0210):

Don't worry eventually you'll learn a trick on how to find the derivative without having to do all this.. work

OpenStudy (anonymous):

I have the correct answer, I just didn't get there the correct way. I avoid the difference quotient whenever possible! :)

OpenStudy (e.mccormick):

See, quotients get no love! They are friendly creatures by their nature, but people see that bar in the middle and think it is dangerous! It starts with the fear of fractions and grows throughout the mathematical career.

OpenStudy (anonymous):

LOL, it's true!